OpenStudy (anonymous):

whats the pattern; 1,8,27,64,125

OpenStudy (anonymous):

cubes

OpenStudy (anonymous):

always in cubes. 1^3 = 1 2^3 = 8 3^3 = 27

OpenStudy (anonymous):

thank you !!!! :)